OpenStudy (anonymous):

Evaluate -9 - (-4). that would be 5 right?

OpenStudy (anonymous):

It would be -4 you distribute that minus sign into the parenthesis (hopefully you know that negative*negative = positive) and you have -9+5= -4

OpenStudy (anonymous):

excuse me I accidentally put -4 when i should have put -5 sorry :p

OpenStudy (anonymous):

Heres a visual -9-(-4) -9+4 -5

OpenStudy (anonymous):

Thank you very much for correcting me and for your explanation.

OpenStudy (anonymous):

No problem have a nice day! Make sure to close the question.

OpenStudy (anonymous):

Evaluate (5 + 2)2 - 40 + 3 = __________. do you mind helping me with this one. I don't understand algebra?

OpenStudy (solomonzelman):

`(1) 2 negative signs become a plus.` \(\large\color{black}{ \displaystyle \color{red}{\rm a}-(-\color{blue}{\rm b})=\color{red}{\rm a}+\color{blue}{\rm b} }\) ---------------------------------------------------------- `(2) When you combine numbers, the sign goes after the greater number.` (for example) \(\large\color{black}{ \displaystyle \color{red}{\rm 6}-\color{blue}{\rm 4}=2}\) or \(\large\color{black}{ \displaystyle -\color{blue}{\rm 4}+\color{red}{\rm 6}=2}\) and on the other hand, \(\large\color{black}{ \displaystyle \color{red}{\rm 4}-\color{blue}{\rm 6}=-2}\) or \(\large\color{black}{ \displaystyle -\color{blue}{\rm 6}+\color{red}{\rm 4}=-2}\) ----------------------------------------------------------

OpenStudy (anonymous):

For (5 + 2)2 - 40 + 3 = you start by working out the parenthesis first. What is the sum in the parenthesis? (5+2)

OpenStudy (anonymous):

7 and 41

OpenStudy (anonymous):

Stay with the parenthesis! So you have (7)2-40 (7)2 is multiplication which always comes before addition and subraction so what is 7*2?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

correct! Now what is 14-40?

OpenStudy (anonymous):

26

OpenStudy (anonymous):

wrong, remember that there is a bigger number being subtracted from a smaller number so its going to be negative

OpenStudy (anonymous):

The answer would be -26

OpenStudy (anonymous):

oh no I forgot to put the negative sign. yes it should be -26.

OpenStudy (anonymous):

There you go. Some other people need some help so have a nice day :D

OpenStudy (anonymous):

you do the same :). thanks again!